STOCHASTIC ADAMS THEOREM FOR A GENERAL COMPACT MANIFOLD

RÉMI LÉANDRE

Abstract. We give a stochastic analoguous of the theorem of Adams, which says that the Hochschild cohomology is equal to the cohomology of the based smooth loop space. The key tools are the stochastic Chen iterated integrals as well as Driver’s flow.